Find the lengths of the line segments joining these pairs of points. Assume that a>0.
(12a,5b) and (3a,5b)

Answer by algebriac(10)

Two points as given in question are (12a,5b) and (3a,5b)
length of a line segment is given by = sqrt( (( X1-X2) ^2) + ((Y1-Y2)^2)) )
Thus
= sqrt( (( 12a-3a) ^2) + ((5b-5b)^2)) )
= sqrt ((9a ^2) +0 )
length= 9a